OpenStudy (anonymous):

i can't find the slope to (3,-4) and (-6,1) can you help please

OpenStudy (anonymous):

u need to find the slope?

OpenStudy (anonymous):

Ya it says to find the slope of the line passing through with the given points (3,-4) (-6,1)

OpenStudy (anonymous):

do u no the slope formula?

OpenStudy (anonymous):

I think it is m=y2-y x2-x1 but Im not sure Im all confused

OpenStudy (anonymous):

I thought it was\[\frac{ x2-x1 }{ y2-y1 }\] im not suree....u should look it up fist

OpenStudy (tkhunny):

Upside-down! Get the4 y's in the numerator.

OpenStudy (anonymous):

Ok but idk how to work it

OpenStudy (tkhunny):

You can't subtract? (3,-4) and (-6,1) y2 - y1 ==> 1 - (-4) = 1 + 4 = 5 x2 - x1 ==> -6 - 3 = -9 I'm pretty sure you can do that.

OpenStudy (anonymous):

I did that and got it wrong so um ya

OpenStudy (anonymous):

This is what you should do: first take the slope formula which is \[m=\frac{ y _{2-} y _{1}}{ {x _{2}-x _{1}} }\] so you plugin the points and you get \[m=\frac{ 1-\left( -4 \right) }{-6-3 }=\frac{ 1+4}{ -9 }=\frac{ 5 }{ -9}=-\frac{ 5 }{ 9}\] Hope this helps! give me a medal if it does :) I'd appreciate it